### v1

# Poincaré inequality on the monotone sets revisited

Abstract
In this short note, we provide a simple inductive proof of the Poincaré inequality on the monotone sets, which is firstly established by Fei and Pinto Jr. [2] using directed heat semigroup method. Such inequality improves the mixing time upper bound of the censored random walk on constant-density monotone sets from $O(n^{3})$ by Ding and Mossel [1] to $O(n^{2})$.
1 Introduction
Discrete-time random walks on finite graphs $G=(V,E)$ model stochastic processes where a walker iteratively transitions between adjacent vertices. Formally, at each timestep $t$, the walker moves from current vertex $u$ to neighbor $v$ with probability $P(u,v)=\frac{1}{{\deg}(u)}$, defining a Markov chain with state space $V$ and transition matrix $P$. The stationary distribution $\pi$, satisfying $\pi P=\pi$, is given by $\pi(v)=\frac{\deg(v)}{2|E|}$ for connected non-bipartite graph. The mixing time $t_{{\rm mix}}(\varepsilon)$ quantifies the convergence rate to stationarity, defined as the minimal $t$ such that:
$\max_{u\in V}\|P^{t}(u,\cdot)-\pi\|_{{\rm TV}}\leq\varepsilon,$ (1)
where $\|\cdot\|_{{\rm TV}}$ denotes the total variation, and $\varepsilon\in(0,1)$ governs approximation accuracy.
An important example is the $n$-dimensional discrete cube. This is a graph $Q_{n}=(\{0,1\}^{n},E)$ with $2^{n}$ vertices indexed by $n$-dimensional Boolean vector. Two 0-1 vectors are adjacent if they differ in exactly one coordinate. We say a set $A\subseteq\{0,1\}^{n}$ is monotone if $x\in A$ and $x\leq y$ imply that $y\in A$. Here, $x\leq y$ denotes $x_{i}\leq y_{i}$ for all $i\in[n]$. Ding and Mossel [1] introduced the study of the random walk on $\{0,1\}^{n}$ censored to $A$. This walk, when at state $x\in A$, attempts to flip a uniformly random coordinate; it moves to the neighbor if it resides in $A$, otherwise it stays at $x$. Note that if we do not add some properties on $A$, the censored random walk may mix well or extremely poorly even when $A$ is large and connected, see Example 1 and 2 in [2]. A significant open question (Question 1.1 in [1]) asked if the optimal mixing time $O_{c}(n\log n)$ holds for $\mu(A)\geq c$, where we use $\mu:=\mu_{1/2}^{\otimes n}$ to denote the uniform probability measure on the cube.
Question 1.1 (Question 1.1 in [1]).
Suppose that there exists a constant $c>0$ such that a monotone subset $A\subseteq\{0,1\}^{n}$ has measure $\mu(A)\geq c$. Consider the censored random walk on $A$ defined above. Is it true that $t_{{\rm mix}}(\varepsilon)\leq O_{\varepsilon,c}(n\log n)$?
When the density $\mu(A)$ is constant, Ding and Mossel established a mixing time bound of $O(n^{3})$. Very recently, Fei and Pinto Jr. [2] show an improved mixing time bound of $O(n^{2})$ for constant-density monotone sets via establishing the following Poincaré inequality on the non-product space.
Theorem 1.2 [2] Theorem 1.6).
Let $A\subseteq\{0,1\}^{n}$ be a non-empty monotone set. We have for all $f:A\to\mathbb{R}$,
${\rm Var}_{A}[f]\leq\frac{1}{1-\sqrt{1-\mu(A)}}\cdot\mathcal{E}_{A}(f).$ (2)
Here ${\rm Var}_{A}[f]$ stands for the variance of $f(x)$ where $x$ is a uniformly random element of $A$ and $\mathcal{E}_{A}(f)$ is called “restricted” Dirichlet form of a function on $A$ defined as follows
$\mathcal{E}_{A}(f)=\mathcal{E}_{A}(f,f):=\frac{1}{4}\cdot\sum\limits_{i=1}^{n}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right],$ (3)
where the expectation is taken with respect to the uniform probability measure on the monotone subset $A\subseteq\{0,1\}^{n}$ and $x^{\oplus i}$ is obtained from $x$ by flipping the $i$th coordinate.
Note that when $A=Q_{n}$, the above result recovers the classical Poincaré inequality on the hypercube. Their proof relies heavily on the connections between the directed heat flow and the directed Poincaré inequality. In this note, we provide a simple inductive proof of their Poincaré inequality, though yielding a slightly worse constant (theirs being 1 versus ours at 2). Nevertheless, we still obtain an $O(n^{2})$ mixing time bound for monotone sets $A$ of constant density via standard Markov chain theory (Theorem 12.4 in [3]). Specifically, we will prove the following theorem, a weaker version of Theorem 1.2.
Theorem 1.3.
Let $A\subseteq\{0,1\}^{n}$ be a non-empty monotone set. We have for all $f:A\to\mathbb{R}$,
${\rm Var}_{A}[f]\leq\frac{2}{1-\sqrt{1-\mu(A)}}\cdot\mathcal{E}_{A}(f).$ (4)
2 Preliminaries and Proofs
In this section, we outline some notation and terminology, and present some simple lemmas for inductive proof of Theorem 1.3.
Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V^{n}$. Let
$V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$
and
$A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},\quad A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}.$
Let $a_{0}:=\mu^{n-1}(A_{0}),a_{1}:=\mu^{n-1}(A_{1})$, and note that
$\begin{split}a&:=\mu^{n}(A)=\frac{1}{2^{n}}\sum_{x\in V^{n}}\mathbbm{1}_{A}(x)=\frac{1}{2}\cdot\frac{1}{2^{n-1}}\left(\sum_{(x^{\prime},1)\in V_{1}}\mathbbm{1}_{A}(x^{\prime},1)+\sum_{(x^{\prime},0)\in V_{0}}\mathbbm{1}_{A}(x^{\prime},0)\right)\\ &=\frac{1}{2}\cdot\frac{1}{2^{n-1}}\left(\sum_{x^{\prime}\in V^{n-1}}\mathbbm{1}_{A_{1}}(x^{\prime})+\sum_{x^{\prime}\in V^{n-1}}\mathbbm{1}_{A_{0}}(x^{\prime})\right)=\frac{1}{2}\mu^{n-1}(A_{1})+\frac{1}{2}\mu^{n-1}(A_{0})=\frac{a_{0}+a_{1}}{2}.\end{split}$ (5)
Define $f_{0}:A_{0}\to\mathbb{R},\ f_{1}:A_{1}\to\mathbb{R}$ by
$f_{0}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},0)\quad\text{and}\quad f_{1}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},1).$
Fact 2.1.
If $A$ is an increasing set, then $A_{0}\subset A_{1}$.
Proof.
For any $x^{\prime}\in A_{0}$, we shall see that $x^{\prime}\in A_{1}$. It is easy to see because $(x^{\prime},0)\leq(x^{\prime},1)$ and $A$ is increasing. ∎
By applying conditional expectation expansion to $n$th variable of restricted Dirichlet forms, we can readily reduce the Dirichlet form of the original function to that of the restricted functions.
Lemma 2.2.
For any $f:A\to\mathbb{R}$, we have
$\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$ (6)
Proof.
We use $x^{\prime}$ to denote the first $(n-1)$-coordinates of $x$. Note that
$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\sum\limits_{i=1}^{n-1}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right]+\frac{1}{4}\cdot\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus n}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1-x_{n})\in A\}}\right]\\ &=\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},1)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},0)\in A\}}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1)\in A\}}\right]\\ &=\frac{a_{1}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{1}\}}\right]+\frac{a_{0}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{0}\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right],\end{split}$
where in the second equality, we observe that
$\begin{split}\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}&\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\right]=\frac{1}{|A|}\cdot\sum\limits_{x^{\prime}\in A_{1}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\\ &=\frac{1}{|A|}\cdot\sum\limits_{x^{\prime}\in A_{0}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}=\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right].\end{split}$
Similarly, we can reduce the variance of the original function to that of the restricted functions.
Lemma 2.3.
For any $f:A\to\mathbb{R}$, we have
${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (7)
Proof.
Note that
$\begin{split}{\rm Var}_{A}[f]=\underset{x\sim A}{\mathbb{E}}[|f-\underset{x\sim A}{\mathbb{E}}f|^{2}]=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\end{split}$
Let $\mu=\underset{x\sim A}{\mathbb{E}}[f(x)],\mu_{1}=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]$ and $\mu_{0}=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]$. Then
$\begin{split}\mu&=\frac{1}{|A|}\sum_{x\in A}f(x)=\frac{1}{|A|}\sum_{x^{\prime}\in A_{1}}f(x^{\prime},1)+\frac{1}{|A|}\sum_{x^{\prime}\in A_{0}}f(x^{\prime},0)=\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]+\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\\ &=\frac{a_{1}}{2a}\mu_{1}+\frac{a_{0}}{2a}\mu_{0}.\end{split}$ (8)
Thus
$\begin{split}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]&=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\frac{a_{1}}{2a}\mu_{1}-\frac{a_{0}}{2a}\mu_{0}\right|^{2}\right]=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}+\mu_{1}-\frac{a_{1}}{2a}\mu_{1}-\frac{a_{0}}{2a}\mu_{0}\right|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}+\frac{a_{0}}{2a}(\mu_{1}-\mu_{0})\right|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}\right|^{2}\right]+2\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}-\mu_{1})\cdot\frac{a_{0}}{2a}(\mu_{1}-\mu_{0})\right]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}\\ &={\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}.\end{split}$
Similarly, we have
$\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]={\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}.$
Thus we have
$\begin{split}{\rm Var}_{A}[f]&=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{1}a_{0}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}^{2}+a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(\mu_{1}-\mu_{0})^{2}\end{split}$
3 An inductive proof of Theorem 1.3
Our goal is to prove that for all $f:A\to\mathbb{R}$,
$\mathcal{E}_{A}(f)\geq c\cdot(1-\sqrt{1-a})\cdot{\rm Var}_{A}[f],$ (9)
where $c>0$ is a universal constant (we can take $c=0.5$). It is easy to check that the theorem holds for $n=1$. Let $n\geq 2$, and suppose the statement of the theorem holds when $n$ is replaced by $n-1$. Using Lemma 2.2 to split the Dirichlet form of the original function to two restricted functions:
$\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$ (10)
By inductive assumption, we have
$\begin{split}&\mathcal{E}_{A_{1}}(f_{1})\geq c\cdot(1-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}],\\ &\mathcal{E}_{A_{0}}(f_{0})\geq c\cdot(1-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}].\end{split}$ (11)
Thus our goal is to show that
$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\geq\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}(c-c\sqrt{1-a}){\rm Var}_{A}[f].\end{split}$ (12)
We also use Lemma 2.3 to split the variance of the original function to two restricted functions:
${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}^{2}+a_{0}a_{1}^{2}}{8a^{3}}\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (13)
Thus it suffices to show that
$\begin{split}&\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(c-c\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (14)
It is equivalent to show that
$\begin{split}&ca_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+ca_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}c\frac{a_{1}a_{0}}{a_{0}+a_{1}}(1-\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (15)
Lemma 3.1.
We may assume both $f_{0},f_{1}$ to be constant on $A_{0}$ and on $A_{1}\setminus A_{0}$.
Proof.
Let $\mu_{1}:=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]$, $\alpha:=\mathbb{E}_{x^{\prime}\sim A_{0}}[f_{1}]$, $\beta:=\mathbb{E}_{x^{\prime}\sim A_{1}\setminus A_{0}}[f_{1}]$ and $\gamma:=\mathbb{E}_{x^{\prime}\sim A_{0}}[f_{0}]$. Note that
$\begin{split}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}f_{1}|^{2}]&=\frac{|A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}+\frac{|A_{1}\setminus A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{1}\setminus A_{0}|}\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}\\ &=\frac{a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}],\end{split}$ (16)
and
$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}(x^{\prime})-\mu_{0}-(f_{1}(x^{\prime})-\mu_{1})+\mu_{0}-\mu_{1}|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}-\mu_{0}|^{2}\right]+\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{1}-\mu_{1}|^{2}\right]+2\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]+(\mu_{0}-\mu_{1})^{2}+2(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (17)
For the sake of simplicity, we denote $C_{1}=\frac{1}{2}\cdot a_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}})$ and $C_{0}=\frac{1}{2}\cdot a_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}})$. Then
$\begin{split}&C_{1}\cdot{\rm Var}_{A_{1}}[f_{1}]+C_{0}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+C_{1}\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]\\ &+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]+\frac{a_{0}}{2}(\mu_{0}-\mu_{1})^{2}+a_{0}(\mu_{0}-\mu_{1})(\mu_{1}-\alpha)\end{split}$ (18)
Now we focus on the term
$\begin{split}&\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &=\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)|\mu_{1}-f_{1}(x^{\prime})|^{2}+\left(C_{0}+\frac{a_{0}}{2}\right)|f_{0}(x^{\prime})-\mu_{0}|^{2}+a_{0}(f_{0}(x^{\prime})-\mu_{0})(\mu_{1}-f_{1}(x^{\prime}))\\ &=\frac{1}{|A_{0}|}\sum_{x^{\prime}\in A_{0}}\vec{b}_{x^{\prime}}^{\top}G_{x^{\prime}}\vec{b}_{x^{\prime}},\end{split}$ (19)
where
$\vec{b}_{x^{\prime}}=\begin{bmatrix}f_{0}(x^{\prime})-\mu_{0}\\ \mu_{1}-f_{1}(x^{\prime})\end{bmatrix},\\ G_{x^{\prime}}=\begin{bmatrix}a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}&\frac{a_{0}}{2}\\ \frac{a_{0}}{2}&a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}.\end{bmatrix}\$
We claim that the matrix $G_{x^{\prime}}$ is positive semidefinite. Note that $G_{x^{\prime}}$ is a symmetric matrix, then we shall use the basic fact that a symmetric matrix $G=\begin{bmatrix}b&c\\ c&d\end{bmatrix}\in\mathbb{R}^{2\times 2}$ is positive semidefinite if and only if $b\geq 0,d\geq 0$ and $bd-c^{2}\geq 0$.
It is easy to see that $a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}\geq 0$ and $a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}\geq 0$ as $a_{0}\leq a\leq a_{1}$. Thus it suffices to verify
$\frac{a_{0}^{2}}{4}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-\frac{a_{0}^{2}}{4}\geq 0.$ (20)
Let $s:=\sqrt{1-a_{0}}$ and $t:=\sqrt{1-a_{1}}$, then $0\leq t\leq s\leq 1$, $21-a=(s^{2}+t^{2})/$ 2 and
$\begin{split}&(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-1\\ &=\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-s+1\right)\cdot\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-t+1\right)-1\\ &=-\frac{1}{2}\cdot(s+t-2)\cdot(\sqrt{2}\sqrt{s^{2}+t^{2}}-s-t)\geq 0,\end{split}$ (21)
where the final one is obviously nonnegative as $s+t-2\leq 0$ and $2(s^{2}+t^{2})\geq(s+t)^{2}$. By the second-order condition of convexity, we know that for any $x^{\prime}\in A_{0}$, the quadratic function $G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))$ is a convex function. Thus by Jensen’s inequality, we have
$\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))]\geq G\left(\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})-\mu_{0}],\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[\mu_{1}-f_{1}(x^{\prime})]\right)=G(0,\mu_{1}-\alpha),$
and
$\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]\geq\left(\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[f_{1}-\mu_{1}]\right)^{2}=(\beta-\mu_{1})^{2}.$
Hence replacing $f_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of (15). ∎
Now let $\alpha$ be the value of $f_{1}$ on $A_{0}$ and $\beta$ be the value of $f_{1}$ on $A_{1}\setminus A_{0}$. Let $\gamma$ be the value of $f_{0}$ on $A_{0}$. Then we compute
$\begin{split}&{\rm Var}_{A_{1}}[f_{1}]=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}^{2}]-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}]^{2}\\ &=\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta^{2})-\left(\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta)\right)^{2}\\ &=\frac{1}{a_{1}}(a_{0}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot\beta^{2})-\frac{1}{a_{1}^{2}}(a_{0}\cdot\alpha+(a_{1}-a_{0})\cdot\beta)^{2}\\ &=\frac{a_{1}a_{0}-a_{0}^{2}}{a_{1}^{2}}\cdot(\alpha-\beta)^{2},\end{split}$ (22)
and
${\rm Var}_{A_{0}}[f_{0}]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}^{2}]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}]^{2}=\frac{a_{0}\cdot 2^{n-1}\cdot\gamma^{2}}{2^{n-1}\cdot a_{0}}-\left(\frac{a_{0}\cdot 2^{n-1}\cdot\gamma}{2^{n-1}\cdot a_{0}}\right)^{2}=0.$ (23)
Thus our goal is to prove that
$\begin{split}&c(\sqrt{1-a}-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})\cdot(\alpha-\beta)^{2}+\frac{a_{1}}{2}\cdot(\gamma-\alpha)^{2}\\ &\geq\frac{c}{a_{1}+a_{0}}\cdot(1-\sqrt{1-a})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}.\end{split}$ (24)
Note that
$\sqrt{1-a}-\sqrt{1-a_{1}}=\frac{1-a-(1-a_{1})}{\sqrt{1-a}+\sqrt{1-a_{1}}}=\frac{a_{1}-\frac{a_{1}+a_{0}}{2}}{\sqrt{1-a}+\sqrt{1-a_{1}}}=\frac{a_{1}-a_{0}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})},$ (25)
and
$1-\sqrt{1-a}=\frac{1-(1-a)}{1+\sqrt{1-a}}=\frac{a_{1}+a_{0}}{2(1+\sqrt{1-a})}.$ (26)
Then it suffice to prove that
$\begin{split}&\frac{c\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})}+\frac{a_{1}}{2}\cdot(\alpha-\gamma)^{2}\geq\frac{c}{2(1+\sqrt{1-a})}\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\\ &\Leftrightarrow c\cdot(1+\sqrt{1-a})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\geq 0.\end{split}$ (27)
Note that
$a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)=(a_{0}-a_{1})\cdot(\alpha-\beta)+a_{1}\cdot(\alpha-\gamma)$
After some simple calculations, the following inequality we need to verify:
$\begin{split}&c\cdot(1-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\cdot(a_{0}-a_{1})\cdot(\alpha-\beta)\cdot(\alpha-\gamma)\geq 0.\end{split}$ (28)
If $\alpha=\gamma$, then it is trivially true. Thus we can assume that $\alpha\neq\gamma$ and let $T:=\frac{\alpha-\beta}{\alpha-\gamma}\cdot(a_{1}-a_{0})$. Then the above inequality transforms to a quadratic inequality in $T$:
$\begin{split}&c\cdot(1-\sqrt{1-a_{1}})\cdot T^{2}+2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot T\\ &a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\geq 0\end{split}$ (29)
Now let us rewrite the Left-hand side of the above inequality as a quadratic in $T$ with coefficients depending only on $a_{1},a_{0}$:
$\begin{split}&A:=c\cdot(1-\sqrt{1-a_{1}})\geq 0\\ &B:=2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\\ &C:=a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}}).\end{split}$
It only needs to verify $\Delta=B^{2}-4AC\leq 0$ with $c=0.5$, which will conclude the proof.
Lemma 3.2.
$B^{2}-4AC\leq 0$.
Proof.
Denote $x=a_{0},y=a_{1}$. Simplifying and rearranging, we need to prove that for any $0\leq x\leq y\leq 1$, we have
$\begin{split}&y\cdot\left(\sqrt{1-\frac{x+y}{2}}+\sqrt{1-y}\right)-2(1-\sqrt{1-y})\cdot\left(1+\sqrt{1-\frac{x+y}{2}}-\frac{y}{2}\right)\leq 0\\ &\Leftrightarrow\left(\sqrt{1-\frac{x+y}{2}}+1\right)\cdot(y-2+2\sqrt{1-y})\leq 0.\end{split}$ (30)
It suffices to show that $(y-2+2\sqrt{1-y})\leq 0$. It is trivially true as
$\begin{split}&y-2+2\sqrt{1-y}\leq 0\\ &\Leftrightarrow 2\sqrt{1-y}\leq 2-y\\ &\Leftrightarrow y^{2}\geq 0.\end{split}$ (31)