### v3

# Functional Inequalities and Random Walks on Increasing Subsets of the Hypercube

Abstract
Motivated by random walks on subsets of the hypercube, we prove two discrete functional inequalities on the hypercube. First, we give a short, elementary proof of the Poincaré inequality on increasing subsets of the cube recently established by Fei and Ferreira Pinto Jr [fei2025spectral], which yields an $O(n^{2})$ upper bound on the mixing time of censored random walks, improving upon previous bounds. Second, adapting Samorodnitsky’s induction method [Samorodnitsky17] to the $p$-biased setting, we establish a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions, which recovers the classic biased edge-isoperimetric inequality for increasing sets and identifies increasing subcubes as the extremizers. This result also admits a probabilistic interpretation in terms of maximizing the mean first exit time of biased random walks.
Keywords: Random walks, mixing time, Poincaré inequality, edge-isoperimetry, $p$-biased hypercube, mean first exit time
MSC codes: 06E30, 05C81, 60C05, 60G50
1 Introduction
Functional inequalities on the discrete hypercube/Hamming graph, such as the Poincaré, log-Sobolev, and edge-isoperimetric inequalities, form a central part of modern discrete analysis. They connect Boolean function analysis, discrete isoperimetric problems, and the spectral theory of Markov chains, providing powerful tools for studying concentration of measure, threshold phenomena, and mixing times of random walks. In the classical uniform product setting on the hypercube $\{0,1\}^{n}$, these inequalities are well understood: sharp constants are known (e.g., the exact log-Sobolev constant [BT2006logsobolev]) and elegant proofs arise via tensorization [gross1975logarithmic], semigroup methods [Ivanisvili2024KKL, ivanisvili2020rademacher], or discrete Fourier analysis [Eldan2025Isoperimetric]. In this paper, we depart from the product setting–by restricting to structured subsets or by working under biased measures, and aim at deriving sharp functional inequalities in these spaces and applying them to characterize the convergence behaviors of relevant random walks.
Beyond the product setting, the study of censored random walks on increasing subsets of the hypercube is motivated by foundational challenges in Markov Chain Monte Carlo (MCMC) for constrained sampling [Jer1998, guruswami2016rapidly, MT2006, LP2017]. A canonical example is sampling uniformly from configurations of a statistical mechanics model (e.g., critical percolation [DM2014]) that satisfy a non-local property such as a left-to-right crossing. While the ambient space is a product measure and trivial to sample from, the constraint set forms a highly correlated, nonproduct subset. Rejection sampling becomes statistically inefficient for rare events, prompting the use of local chains that remain within the constraint set. The natural censored dynamics re-randomizes one coordinate at a time and accepts the move only if the property is preserved. This process evolves on an increasing subset, and monotonicity ensures irreducibility. The central algorithmic question is to quantify the mixing time and the related survival/exit characteristics of this censored walk, since these quantities determine the efficiency and robustness of constrained sampling in combinatorics, statistical physics, and property testing.
1.1 Poincaré inequality and random walks on increasing sets
An illustration of non-product structure is motivated from censored random walks, where product structure is explicitly broken. If we censor the simple random walk on $\{0,1\}^{n}$ to a subset $A$ (only moves that stay in $A$ are accepted), the chain now lives on a non-product space with geometry dictated by the boundary of $A$. Currently, there is a lack of good estimation for the mixing time of such random walks. Let us introduce this problem in details.
We work on the $n$-dimensional hypercube. The Hamming graph is the graph $Q^{n}=(\{0,1\}^{n},E)$ with $2^{n}$ vertices indexed by binary strings, where two vertices are adjacent (written $x\sim y$) iff they differ in exactly one coordinate. Discrete-time random walks on the Hamming graph (or more generally, on finite graphs $G=(V,E)$) model stochastic processes where a walker iteratively transitions between adjacent vertices. Formally, at each timestep $t$, the walker moves from current vertex $u$ to neighbor $v$ with probability $P(u,v)=\frac{1}{{\deg}(u)}$, defining a Markov chain with state space $V$ and transition matrix $P$. The stationary distribution $\pi$, satisfying $\pi P=\pi$, is given by $\pi(v)=\frac{\deg(v)}{2|E|}$ for connected non-bipartite graph. The mixing time $t_{{\rm mix}}(\varepsilon)$ quantifies the convergence rate to stationarity, defined as the minimal $t$ such that:
$\max_{u\in V}\|P^{t}(u,\cdot)-\pi\|_{{\rm TV}}\leq\varepsilon,$ (1)
where $\|\cdot\|_{{\rm TV}}$ denotes the total variation, and $\varepsilon\in(0,1)$ governs approximation accuracy.
Ding and Mossel [DM2014] introduced the study of the random walk on $\{0,1\}^{n}$ censored to $A$. This walk, when at state $x\in A$, attempts to flip a uniformly random coordinate; it moves to the neighbor if it resides in $A$, otherwise it stays at $x$. Note that if we do not impose any restrictions on $A$, the censored random walk may mix well, and also may mix extremely poorly even when $A$ is large and connected, see [fei2025spectral, Example 1,2]. A significant open question ([DM2014, Question 1.1]) asked if the optimal mixing time belongs to $O_{c}(n\log n)$ holds for increasing subsets $A$ with constant density under the uniform measure.
Question 1 ([DM2014, Question 1.1]).
Suppose that there exists a constant $c>0$ such that a increasing subset $A\subseteq\{0,1\}^{n}$ has measure $\mu(A)\geq c$. Consider the censored random walk on $A$ defined above. Is it true that $t_{{\rm mix}}(\varepsilon)\leq O_{\varepsilon,c}(n\log n)$?
When the density $\mu(A)$ is constant, Ding and Mossel established a mixing time bound of $O(n^{3})$. Very recently, Fei and Ferreira Pinto Jr [fei2025spectral] make a breakthrough on this topic by showing an improved mixing time bound of $O(n^{2})$ for constant-density increasing sets. Their result is obtained via establishing the following Poincaré inequality on the non-product space.
Theorem 1.1 ([fei2025spectral, Theorem 1.6]).
Let $A\subseteq\{0,1\}^{n}$ be a non-empty increasing set. It holds that for all $f:A\to\mathbb{R}$,
${\rm Var}_{A}[f]\leq\frac{1}{1-\sqrt{1-\mu(A)}}\cdot\mathcal{E}_{A}(f).$ (2)
Here ${\rm Var}_{A}[f]$ stands for the variance of $f(x)$ where $x$ is a uniformly random element of $A$, and $\mathcal{E}_{A}(f)$ is called “restricted” Dirichlet form of a function on $A$ defined as follows
$\mathcal{E}_{A}(f)=\mathcal{E}_{A}(f,f):=\frac{1}{4}\cdot\sum\limits_{i=1}^{n}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right],$ (3)
where the expectation is taken with respect to the uniform probability measure on the increasing subset $A\subseteq\{0,1\}^{n}$ and $x^{\oplus i}$ is obtained from $x$ by flipping the $i$th coordinate.
Note that when $A=Q^{n}$, the above result recovers the classical Poincaré inequality on the hypercube. Their proof relies heavily on the connections between the directed heat flow and the directed Poincaré inequality. In this paper, we provide a simple inductive proof of their Poincaré inequality, though yielding a slightly worse constant (theirs being 1 versus ours at 2). Specifically, we use the induction method to prove the following theorem, a weaker version of Theorem 1.1.
Theorem 1.2.
Let $A\subseteq\{0,1\}^{n}$ be a non-empty increasing set. We have for all $f:A\to\mathbb{R}$,
${\rm Var}_{A}[f]\leq\frac{2}{1-\sqrt{1-\mu(A)}}\cdot\mathcal{E}_{A}(f).$ (4)
Nevertheless, we still obtain an $O(n^{2})$ mixing time bound for increasing sets $A$ of constant density via standard Markov chain theory ([LP2017, Theorem 12.4]).
Theorem 1.3 ([fei2025spectral]).
Let $A\subseteq\{0,1\}^{n}$ be a non-empty increasing set. Then the random walk on $\{0,1\}^{n}$ censored to $A$ has mixing time
$t_{{\rm mix}}\leq\frac{2n}{\mu(A)}\cdot\log(4\cdot 2^{n}\mu(A)).$
Fei and Ferreira Pinto Jr. [fei2025spectral] develop an $L^{2}$ functional–analytic framework for increasing sets $A\subseteq\{0,1\}^{n}$. They introduce a directed gradient that counts only order-consistent edges (down moves), which defines a directed Laplacian $L_{A}^{\to}$ and energy $\mathcal{E}_{A}^{\to}(f)$. They then run the monotone heat flow generated by $L_{A}^{\to}$ and prove an energy–variance dissipation inequality along the flow, yielding a directed Poincaré inequality with constant $C(\mu(A))=O(1)$ whenever $\mu(A)$ is bounded away from 0. A second ingredient is an approximate FKG inequality on increasing sets that supplies quantitative positive-correlation bounds. Together these tools give an optimal (up to constants) spectral gap for the censored walk on $A$.
In contrast, our method is designed to be simple and easy to check. We establish a slightly weaker, yet asymptotically optimal, Poincaré inequality by a classical induction-on-dimension argument. The proof restricts functions to lower-dimensional subcubes and uses the inherent increasing structure $A_{0}\subseteq A_{1}$ to decompose the variance and the Dirichlet form fiberwise, ultimately reducing the task to an explicit five-point inequality that can be verified directly. The resulting constant is within a factor $2$ of the optimal value and already implies the same mixing bound $O(n^{2})$ for censored walks. Because every step is combinatorial and closed-form, the proof is short, self-contained, and readily adaptable.
1.2 Biased Samorodnitsky’s inequality and exit time of random walks from increasing sets
Another illustration of non-product structure is motivated from characterizing the mean first exit time of a random walk from an increasing set.
Let $g:\{0,1\}^{n}\to\mathbb{R}$, we define the Dirichlet form of $g$ as
$\mathcal{E}(g,g)=\underset{x\sim\mu}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}(g(x)-g(y))^{2}\right],$
where the expectation is taken with respect to the uniform probability measure $\mu:=\mu_{1/2}^{\otimes n}$ on the hypercube. Furthermore, we denote by $x\sim A$ that $x$ is drawn uniformly at random from the subset $A\subseteq\{0,1\}^{n}$. In 2017, Samorodnitsky [Samorodnitsky17] established the following discrete functional inequality on the hypercube.
Theorem 1.4 (Samorodnitsky [Samorodnitsky17]).
Let $A$ be a subset of $\{0,1\}^{n}$ and let $g:\{0,1\}^{n}\to\mathbb{R}$ supported on $A$. Then
$\mathcal{E}(g,g)\geq\frac{|A|}{2^{n-1}}\left(\underset{x\sim A}{\mathbb{E}}[|g(x)|]\right)^{2}\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right),$ (5)
equality holds if and only if $g$ is an indicator function of a co-dimension $k$ subcube for any $0\leq k\leq n$.
When $g$ is an indicator function of $A$, Samorodnitsky’s inequality Eq. 5 implies the classic edge-isoperimetric inequality on $Q^{n}$: for every $A\subset\{0,1\}^{n}$,
$|\nabla A|\geq|A|\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right),$ (6)
where the edge boundary $\nabla(A)$ is the set of edges between $A$ and its complement $A^{c}=V(Q^{n})\setminus A$. Equality in Eq. 6 holds if and only if $A$ is a subcube of co-dimension $k$, for any $0\leq k\leq n$.
Besides the motivation from generalization of Haper’s edge-isoperimetric inequality, Samorodnitsky’s inequality admits a natural interpretation in languages of random walk on the hypercube. Specifically, it is equivalent to a sharp estimate of mean first exit time of a random walk from a subset with a given size [Samorodnitsky17].
We now generalize Samorodnitsky’s inequality to the $p$-biased setting so that the resulting functional inequality recovers the sharp constant in the $p$-biased edge-isoperimetric inequality. Let $f\in L^{2}(\{0,1\}^{n},\mu_{p}^{\otimes n})$, where $\mu_{p}$ is the Bernoulli measure on $\{0,1\}$ with $\mu_{p}(1)=p$ and $\mu_{p}(0)=1-p$. For $x\in\{0,1\}^{n}$, write $|x|=|\{i\in[n]:x_{i}=1\}|$, so that
$\mu_{p}(x):=\mu_{p}^{\otimes n}(x)=p^{|x|}(1-p)^{n-|x|},\quad\mathbb{E}_{p}[f]:=\underset{x\sim\mu_{p}}{\mathbb{E}}[f(x)]=\sum\limits_{x\in\{0,1\}^{n}}f(x)\mu_{p}(x).$
Many applications (e.g., to the study of percolation [BKS1999], threshold phenomena in random graphs [KLMM2024globalhyper2024, Talagrand1994randomgraph], and hardness of approximation [DS2005Annal]) rely upon the use of the biased measure on the discrete cube. The definition of Dirichlet quadratic form of $g$ with respect to the biased measure is
$\mathcal{E}_{p}(g,g):=\underset{x\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}(g(x)-g(y))^{2}\right].$
We consider increasing functions defined on the discrete cube $\{0,1\}^{n}$ endowed with the $p$-biased measure.
Definition 1.
A function $g:(\{0,1\}^{n},\mu_{p})\to\mathbb{R}$ is increasing if for all $x,y\in\{0,1\}^{n}$,
$x_{i}\leq y_{i},\ \forall i\in[n]\Rightarrow g(x)\leq g(y).$
A subset $A\subset\{0,1\}^{n}$ is called increasing if its characteristic function is increasing.
The following ‘biased’ generalization of Harper’s edge-isoperimetric inequality for increasing subsets is considered folklore (see [EKL2019biased, Kahnkalai07]).
Theorem 1.5.
Let $A$ be an increasing subset of $\{0,1\}^{n}$ and let $0<p<1$. Then
$p\cdot\mathcal{E}_{p}(\mathbbm{1}_{A},\mathbbm{1}_{A})\geq\mu_{p}(A)\log_{p}\mu_{p}(A),$ (7)
equality holds in Eq. 7 if and only if $A$ is an increasing subcube.
Inspired by Eq. 5, it is natural to seek a functional inequality on the $p$-biased cube for general real-valued (increasing) functions that, in the special case of indicator functions, reduces to the $p$-biased edge-isoperimetric inequality with the sharp constant. We establishes such an inequality, which recovers Eq. 7 by substituting $g=\mathbbm{1}_{A}$.
Theorem 1.6.
Let $A$ be a subset of $\{0,1\}^{n}$ and let $g$ be a real-valued increasing function on $\{0,1\}^{n}$ with support $A$. Then for any $0<p<1$,
$p\cdot\mathcal{E}_{p}(g,g)\geq\frac{\mathbb{E}_{p}[|g|]^{2}}{\mu_{p}(A)}\log_{p}\mu_{p}(A),$ (8)
equality holds if and only if $g$ is an indicator function of an increasing subcube of co-dimension $k$ for any $0\leq k\leq n$.
We view the $p$-biased hypercube as a weighted undirected graph $Q^{n}_{p}$, where for every vertex $x\in\{0,1\}^{n}$ and coordinate index $i\in[n]$, the edge weight is defined by
$\begin{split}w_{p}(\{x,x\oplus e_{i}\})&=\mu_{p}^{\otimes n}(x)+\mu_{p}^{\otimes n}(x\oplus e_{i})=\mu_{p}^{\otimes(n-1)}(x_{\backslash i}),\end{split}$
where $x_{\backslash i}=(x_{1},...,x_{i-1},x_{i+1},...,x_{n})$. Thus each edge is weighted proportionally to the combined $p$-biased mass of its two endpoints.
A random walk on a graph is a process that begins at some vertex, and at each time step moves to another vertex. When the graph is unweighted (i.e., $p=1/2$), the vertex the (simple) random walk moves to is chosen uniformly at random among the neighbors of the present vertex. When the graph is weighted, it moves to a neighbor with probability proportional to the weight of the corresponding edge. In [Samorodnitsky17], Samorodnitsky observed that the functional inequality Eq. 5 is equivalent to a probabilistic statement about random walks on the hypercube: subcubes maximize mean first exit time among all subsets of the same cardinality. Using the same argument, our new $p$-biased edge-isoperimetric inequality Eq. 8 admits an analogous probabilistic interpretation in terms of a simply random walk on the weighted hypercube $Q^{n}_{p}$.
Let $A\subseteq\{0,1\}^{n}$ and define the following random walk $X_{t}$. The random walk starts at a random point $X_{0}=x\in A$ following from the $p$-biased conditional law, i.e., $\nu(x):=\frac{\mu_{p}(x)}{\mu_{p}(A)},\forall x\in A$, and at each step, moves to one of its $n$ neighbors with probability ${\rm P}_{p}(x,x\oplus e_{i})=\omega_{p}(\{x,x\oplus e_{i}\})/d_{p}(x)$, where $d_{p}(x)=\sum_{i=1}^{n}\omega_{p}(\{x,x\oplus e_{i}\})$ is the weighted degree of vertex $x$. Let $Y$ be the (random) time that it takes the walk to exit $A$ for the first time for the $p$-biased walk. We refer to $T(A,p)=\mathbb{E}[Y]$ as the mean first exit time of $A$ under the $p$-biased walk. We further introduce the weighted mean first exit time as
$\tilde{T}(A,p):=\mathbb{E}\left[\sum\limits_{t=0}^{Y-1}\phi(X_{t})\right],{\rm where}\ \phi(x):=\frac{\mu_{p}(x)}{d_{p}(x)}.$
Theorem 1.7.
Let $A\subset\{0,1\}^{n}$ be increasing and $0<p<1$. Then
  • (1)
    $\tilde{T}(A,p)\leq\frac{p}{\log_{p}\mu_{p}(A)}$ with equality if and only if $A$ is an increasing subcube.
  • (2)
    $T(A,p)\leq\frac{p}{\min\{p,1-p\}}\cdot\frac{n}{\log_{p}(\mu_{p}(A))}$ with equality if and only if $p=\tfrac{1}{2}$ and $A$ is an increasing subcube.
Example 1 (Mean first exit time for the $p$-biased walk on increasing subcubes).
Fix $0<p<1$ and $n\in\mathbb{N}$. Let $C:=\{x\in\{0,1\}^{n}:x_{1}=\cdots=x_{k}=1\}$ be the increasing subcube of codimension $k$, i.e., $k=\log_{p}\mu_{p}(A)$. Then
$T(C,p)=1+2p\left(\frac{n}{k}-1\right).$
In particular, for $p=\tfrac{1}{2}$ this reduces to $T(C,p)=n/k$.
1.3 Organization
This paper is organized as follows. Section 2 presents our first main result: an inductive proof of a Poincaré inequality on increasing subsets of the cube, which yields an $O(n^{2})$ upper bound on the mixing time of the censored random walk. The arguments follow the induction + five-point paradigm, though the analytic verification is much simple. Section 3 turns to our second main result: a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions. We develop the induction-on-dimension framework and verify the required two-point inequality in this setting. Section 4 explains how inequality Eq. 8 in Theorem 1.6 leads to bounds on the mean survival (first-exit) time of random walks from increasing sets (Theorem 1.7). Section 5 verifies the mean survival time of random walks from increasing subcubes (1). Finally, Section 6 highlight a direction for future work: establishing suitable (possibly modified) log-Sobolev inequalities on increasing sets, which are closely related to the Ding–Mossel conjecture.
2 Proof of Theorem 1.2
2.1 Preliminaries
In this subsection, we outline some notation and terminology, and present some simple lemmas for inductive proof of Theorem 1.2. In the language of Boolean functions, there are two different ways to prove inequalities about $f:\{0,1\}^{n}\to\mathbb{R}$ by induction on $n$. One method, which might be called “induction by derivatives”, decomposes a function $f$ as follows: $f(x)=(-1)^{x_{i}}g_{1}(y)+g_{0}(y)$ for some functions $g_{0},g_{1}$ on the (n1)(n-1)-dimensional cube, where $y=(x_{1},\dots,x_{n-1})$. Such a decomposition is motivated from Fourier expansion. This induction method underlies classic proofs of Bonami’s lemma and the hypercontractive inequality on the hypercube [Ryanbook2014]. The other method, which we term “induction by restrictions”, proceeds via the restricted functions $g_{a}$ obtained by fixing the last coordinate to $a\in\{0,1\}$. One then expresses Dirichlet forms (or total influence) and variances recursively in terms of $g_{0}$ and $g_{1}$, applies the inductive hypothesis in dimension n1n-1, and controls the remaining cross-terms by verifying an explicit two-point (and occasionally four/five-point) inequality. We prove both results within the standard induction-by-restrictions, widely used in combinatorics and discrete analysis (see, e.g., [alpay2025lowerboundsdyadicsquare, beltran2023sharpisoperimetricinequalitieshypercube, BIMP2025young, CTKS2025inequalities, durcik2024sharpisoperimetricinequalitieshamming, EKL2019biased, FS2007CPCKKL, Kahnkalai07, Samorodnitsky17, yu2025averagedistancelevel1fourier]).
Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V^{n}$. Let $V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$ and $A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}$.
Let $a_{0}:=\mu^{n-1}(A_{0}),a_{1}:=\mu^{n-1}(A_{1})$, and note that
$\begin{split}a:=\mu^{n}(A)=\frac{1}{2}\mu^{n-1}(A_{1})+\frac{1}{2}\mu^{n-1}(A_{0})=\frac{a_{0}+a_{1}}{2}.\end{split}$ (9)
Define $f_{0}:A_{0}\to\mathbb{R},\ f_{1}:A_{1}\to\mathbb{R}$ by
$f_{0}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},0)\quad\text{and}\quad f_{1}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},1).$
By expanding the Dirichlet form with respect to the conditional expectation on the $n$th variable, we can readily express the Dirichlet form of the original function in terms of those of its restricted versions.
Lemma 2.1.
For any $f:A\to\mathbb{R}$, we have
$\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$ (10)
Proof.
We use $x^{\prime}$ to denote the first (n1)(n-1)-coordinates of $x$. First, separate the contribution of the last coordinate from that of the first n1n-1 coordinates; this gives
$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\sum\limits_{i=1}^{n-1}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right]\\ &+\frac{1}{4}\cdot\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus n}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1-x_{n})\in A\}}\right]\end{split}$
Next, write $x=(x^{\prime},x_{n})$ with $x^{\prime}\in\{0,1\}^{n-1}$ and condition on $x_{n}\in\{0,1\}$. Then
$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},1)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1)\in A\}}\right]\end{split}$
Now split according to whether the flipped point stays in the same fiber $A_{b}$ for $b\in\{0,1\}$ or moves across fibers; this becomes
$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{1}\}}\right]\\ &+\frac{a_{0}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{0}\}}\right]\\ &+\frac{|A_{0}|}{4|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right]+\frac{|A_{0}|}{4|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right],\end{split}$
where in the second equality, we observe that
$\begin{split}\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}&\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\right]=\frac{1}{|A|}\sum\limits_{x^{\prime}\in A_{1}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\\ &=\frac{1}{|A|}\sum\limits_{x^{\prime}\in A_{0}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}=\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right].\end{split}$
Similarly, by conditioning on the $n$th coordinate, the variance of the original function can be expressed in terms of the variances of its restricted functions.
Lemma 2.2.
For any $f:A\to\mathbb{R}$, we have
$\begin{split}{\rm Var}_{A}[f]=&\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]\\ &+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.\end{split}$ (11)
Proof.
Note that
$\begin{split}{\rm Var}_{A}[f]=\underset{x\sim A}{\mathbb{E}}[|f-\underset{x\sim A}{\mathbb{E}}f|^{2}]=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}].\end{split}$
Let $\mu=\underset{x\sim A}{\mathbb{E}}[f(x)],\mu_{1}=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]$ and $\mu_{0}=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]$. Then
$\mu=\frac{1}{|A|}\sum_{x\in A}f(x)=\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]+\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]=\frac{a_{1}}{2a}\mu_{1}+\frac{a_{0}}{2a}\mu_{0}.$ (12)
Thus by the formula $\mathbb{E}[(f-b)^{2}]={\rm Var}[f]+(\mathbb{E}[f]-b)^{2}$, we have that
$\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]={\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2},$
and
$\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]={\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}.$
Thus we have
$\begin{split}{\rm Var}_{A}[f]&=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{1}a_{0}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(\mu_{1}-\mu_{0})^{2}\end{split}$
2.2 Induction part
Our goal is to prove that for every function $f:A\to\mathbb{R}$,
$\mathcal{E}_{A}(f)\geq c\cdot(1-\sqrt{1-a})\cdot{\rm Var}_{A}[f],$ (13)
where $c>0$ is a universal constant (we may take $c=0.5$). It is easy to verify that the theorem holds for $n=1$. Let $n\geq 2$, and suppose the statement of the theorem holds when $n$ is replaced by n1n-1. Applying Lemma 2.1 to split the Dirichlet form of the original function into two restricted functions:
$\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$
By inductive assumption, we have
$\begin{split}&\mathcal{E}_{A_{1}}(f_{1})\geq c\cdot(1-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}],\\ &\mathcal{E}_{A_{0}}(f_{0})\geq c\cdot(1-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}].\end{split}$ (14)
Thus our goal is to show that
$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\geq\frac{a_{1}}{2a}(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]\\ &+\frac{a_{0}}{4a}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}(c-c\sqrt{1-a}){\rm Var}_{A}[f].\end{split}$ (15)
We also use Lemma 2.2 to split the variance of the original function to two restricted functions:
${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$
Thus it suffices to show that
$\begin{split}&\frac{a_{1}(c-c\sqrt{1-a_{1}})}{2a}{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}(c-c\sqrt{1-a_{0}})}{2a}{\rm Var}_{A_{0}}[f_{0}]\\ &\quad\ +\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}\frac{a_{1}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{0}}[f_{0}]\\ &\quad\ +\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}(c-c\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (16)
It is equivalent to show that
$\begin{split}&ca_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+ca_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]\\ &+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\underset{?}{\geq}c\frac{a_{1}a_{0}}{a_{0}+a_{1}}(1-\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (17)
Lemma 2.3.
We may assume both $f_{0},f_{1}$ to be constant on $A_{0}$ and on $A_{1}\setminus A_{0}$.
Proof.
Let $\mu_{1}:=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]$, $\alpha:=\mathbb{E}_{x^{\prime}\sim A_{0}}[f_{1}]$, $\beta:=\mathbb{E}_{x^{\prime}\sim A_{1}\setminus A_{0}}[f_{1}]$ and $\gamma:=\mathbb{E}_{x^{\prime}\sim A_{0}}[f_{0}]$. Note that
$\begin{split}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}f_{1}|^{2}]&=\frac{|A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}\\ &+\frac{|A_{1}\setminus A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{1}\setminus A_{0}|}\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}\\ &=\frac{a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}],\end{split}$ (18)
and
$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}(x^{\prime})-\mu_{0}-(f_{1}(x^{\prime})-\mu_{1})+\mu_{0}-\mu_{1}|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}-\mu_{0}|^{2}\right]+\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{1}-\mu_{1}|^{2}\right]+2\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &+(\mu_{0}-\mu_{1})^{2}+2(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (19)
For the sake of simplicity, we denote $C_{1}=\frac{1}{2}\cdot a_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}})$ and $C_{0}=\frac{1}{2}\cdot a_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}})$. Then
$\begin{split}&C_{1}\cdot{\rm Var}_{A_{1}}[f_{1}]+C_{0}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+C_{1}\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]\\ &+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &+\frac{a_{0}}{2}(\mu_{0}-\mu_{1})^{2}+a_{0}(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (20)
Now we focus on the term
$\begin{split}&\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]\\ &\quad+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &=\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)|\mu_{1}-f_{1}(x^{\prime})|^{2}+\left(C_{0}+\frac{a_{0}}{2}\right)|f_{0}(x^{\prime})-\mu_{0}|^{2}\\ &\quad+a_{0}(f_{0}(x^{\prime})-\mu_{0})(\mu_{1}-f_{1}(x^{\prime}))\\ &=\frac{1}{|A_{0}|}\sum_{x^{\prime}\in A_{0}}\vec{b}_{x^{\prime}}^{\top}G_{x^{\prime}}\vec{b}_{x^{\prime}},\end{split}$ (21)
where
$\vec{b}_{x^{\prime}}=\begin{bmatrix}f_{0}(x^{\prime})-\mu_{0}\\ \mu_{1}-f_{1}(x^{\prime})\end{bmatrix},\\ G_{x^{\prime}}=\begin{bmatrix}a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}&\frac{a_{0}}{2}\\ \frac{a_{0}}{2}&a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}\end{bmatrix}$.\\
We claim that the matrix $G_{x^{\prime}}$ is positive semidefinite. Note that $G_{x^{\prime}}$ is a symmetric matrix, then we shall use the basic fact that a symmetric matrix $G=\begin{bmatrix}b&c\\ c&d\end{bmatrix}\in\mathbb{R}^{2\times 2}$ is positive semidefinite if and only if $b\geq 0,d\geq 0$ and $bd-c^{2}\geq 0$.
It is easy to see that $a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}\geq 0$ and $a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}\geq 0$ as $a_{0}\leq a\leq a_{1}$. Thus it suffices to verify
$\frac{a_{0}^{2}}{4}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-\frac{a_{0}^{2}}{4}\geq 0.$ (22)
Let $s:=\sqrt{1-a_{0}}$ and $t:=\sqrt{1-a_{1}}$, then $0\leq t\leq s\leq 1$, $1-a=(s^{2}+t^{2})/2$ and
$\begin{split}&(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-1\\ &=\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-s+1\right)\cdot\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-t+1\right)-1\\ &=-\frac{1}{2}\cdot(s+t-2)\cdot(\sqrt{2}\sqrt{s^{2}+t^{2}}-s-t)\geq 0,\end{split}$ (23)
where the final one is obviously nonnegative as $s+t-2\leq 0$ and $2(s^{2}+t^{2})\geq(s+t)^{2}$. By the second-order condition of convexity, we know that for any $x^{\prime}\in A_{0}$, the quadratic function $G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))$ is a convex function. Thus by Jensen’s inequality, we have
$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))]\\ &\geq G\left(\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})-\mu_{0}],\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[\mu_{1}-f_{1}(x^{\prime})]\right)=G(0,\mu_{1}-\alpha),\end{split}$
and
$\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]\geq\left(\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[f_{1}-\mu_{1}]\right)^{2}=(\beta-\mu_{1})^{2}.$
Hence replacing $f_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of Eq. 17. ∎
2.3 Five-point Inequality Part
Now let $\alpha$ be the value of $f_{1}$ on $A_{0}$ and $\beta$ be the value of $f_{1}$ on $A_{1}\setminus A_{0}$. Let $\gamma$ be the value of $f_{0}$ on $A_{0}$. Then we compute
$\begin{split}&{\rm Var}_{A_{1}}[f_{1}]=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}^{2}]-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}]^{2}\\ &=\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta^{2})\\ &-\left(\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta)\right)^{2}\\ &=\frac{1}{a_{1}}(a_{0}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot\beta^{2})-\frac{1}{a_{1}^{2}}(a_{0}\cdot\alpha+(a_{1}-a_{0})\cdot\beta)^{2}\\ &=\frac{a_{1}a_{0}-a_{0}^{2}}{a_{1}^{2}}\cdot(\alpha-\beta)^{2},\end{split}$
and
${\rm Var}_{A_{0}}[f_{0}]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}^{2}]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}]^{2}=\frac{a_{0}\cdot 2^{n-1}\cdot\gamma^{2}}{2^{n-1}\cdot a_{0}}-\left(\frac{a_{0}\cdot 2^{n-1}\cdot\gamma}{2^{n-1}\cdot a_{0}}\right)^{2}=0.$
Thus to verify Eq. 17, it suffices to prove the following five-point inequality with $c=\frac{1}{2}$.
$\begin{split}&c(\sqrt{1-a}-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})\cdot(\alpha-\beta)^{2}+\frac{a_{1}}{2}\cdot(\gamma-\alpha)^{2}\\ &\geq\frac{c}{a_{1}+a_{0}}\cdot(1-\sqrt{1-a})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}.\end{split}$ (24)
Note that
$\sqrt{1-a}-\sqrt{1-a_{1}}=\frac{1-a-(1-a_{1})}{\sqrt{1-a}+\sqrt{1-a_{1}}}=\frac{a_{1}-a_{0}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})},$
and
$1-\sqrt{1-a}=\frac{1-(1-a)}{1+\sqrt{1-a}}=\frac{a_{1}+a_{0}}{2(1+\sqrt{1-a})}.$
Then it suffice to prove that
$\begin{split}&\frac{c\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})}+\frac{a_{1}}{2}\cdot(\alpha-\gamma)^{2}\\ &\geq\frac{c}{2(1+\sqrt{1-a})}\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\\ &\Longleftrightarrow c\cdot(1+\sqrt{1-a})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}\\ &\quad\quad\ +a_{1}\cdot(1+\sqrt{1-a})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &\quad\quad\ -c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\geq 0.\end{split}$ (25)
Note that $a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)=(a_{0}-a_{1})\cdot(\alpha-\beta)+a_{1}\cdot(\alpha-\gamma).$ After some simple calculations, the inequality to be verified becomes
$\begin{split}&c(1-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}\\ &+a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\cdot(a_{0}-a_{1})\cdot(\alpha-\beta)\cdot(\alpha-\gamma)\geq 0.\end{split}$ (26)
If $\alpha=\gamma$, the inequality is trivially satisfied. Thus, assume $\alpha\neq\gamma$ and set $T:=\frac{\alpha-\beta}{\alpha-\gamma}\cdot(a_{1}-a_{0})$. Then the above inequality can be rewritten as a quadratic in $T$:
$\begin{split}&c\cdot(1-\sqrt{1-a_{1}})\cdot T^{2}+2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot T\\ &a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\geq 0.\end{split}$ (27)
We may rewrite the left-hand side as a quadratic polynomial in $T$ ($AT^{2}+BT+C$), whose coefficients depend only on $a_{1},a_{0}$:
$\begin{split}&A:=c\cdot(1-\sqrt{1-a_{1}})\geq 0\\ &B:=2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\\ &C:=a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}}).\end{split}$
To ensure the quadratic is nonnegative for all $T$, it suffices to check that its discriminant satisfies $\Delta=B^{2}-4AC\leq 0$ with $c=0.5$, which will conclude the proof.
Lemma 2.4.
$B^{2}-4AC\leq 0$.
Proof.
Denote $x=a_{0},y=a_{1}$. Simplifying and rearranging, we need to prove that for any $0\leq x\leq y\leq 1$, we have
$\begin{split}&y\cdot\left(\sqrt{1-\frac{x+y}{2}}+\sqrt{1-y}\right)-2(1-\sqrt{1-y})\cdot\left(1+\sqrt{1-\frac{x+y}{2}}-\frac{y}{2}\right)\leq 0\\ &\Leftrightarrow\left(\sqrt{1-\frac{x+y}{2}}+1\right)\cdot(y-2+2\sqrt{1-y})\leq 0.\end{split}$
It suffices to show that $(y-2+2\sqrt{1-y})\leq 0$. It is trivially true as
$\begin{split}&y-2+2\sqrt{1-y}\leq 0\\ &\Leftrightarrow 2\sqrt{1-y}\leq 2-y\\ &\Leftrightarrow y^{2}\geq 0.\end{split}$
3 Proof of Theorem 1.6
3.1 Preliminaries
In this subsection, we outline some notation and terminology, and present some simple lemmas for inductive proof of Theorem 1.6.
For a parameter $p\in(0,1)$, the $p$-biased measure on $\{0,1\}$ assigns probability $p$ to $1$ and probability 1p1-p to 0. Let $\mu_{p}^{\otimes n}$ be the product distribution of $\mu_{p}$ over $\{0,1\}^{n}$. For $x\in\{0,1\}^{n}$, let $|x|=|\{i\in[n]:x_{i}=1\}|$ and then
$\mu_{p}(x):=\mu_{p}^{\otimes n}(x)=p^{|x|}(1-p)^{n-|x|}.$
We consider the space of real valued functions on $\{0,1\}^{n}$, equipped with the inner product
$\left\langle f,g\right\rangle_{p}=\underset{x\sim\mu_{p}}{\mathbb{E}}[f(x)g(x)]=\int_{\{0,1\}^{n}}f(x)g(x)\mu_{p}(x)=\sum\limits_{x\in\{0,1\}^{n}}f(x)\cdot g(x)\mu_{p}(x).$
Definition 2.
For $q\geq 1$, the $q$-norm w.r.t. $p$-biased measure of $f:\{0,1\}^{n}\to\mathbb{R}$ is
$\|f\|_{q,\mu_{p}}:=\left(\underset{x\sim\mu_{p}}{\mathbb{E}}[|f(x)|^{q}]\right)^{1/q}.$
Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V$. Let $V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$ and $A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}.$
Let $\mu_{p}^{n-1}(A_{0})=a_{0},\mu_{p}^{n-1}(A_{1})=a_{1}$, and $\mu_{p}^{n}(A)=a=pa_{1}+(1-p)a_{0}$. Define $g_{0},g_{1}:(\{0,1\}^{n-1},\mu_{p}^{n-1})\to\mathbb{R}$ by $g_{0}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},0)$ and $g_{1}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},1)$ and the support of $g_{1}$ and $g_{0}$ are the whole set $A_{1}$ and $A_{0}$. Note that if $g$ is an increasing function, then the set $A$ is also an increasing set, since it is the support of an increasing function $g$. This implies $A_{0}\subset A_{1}$.
Fact 3.1.
If $A$ is an increasing set, then $A_{0}\subset A_{1}$.
Proof.
For any $x^{\prime}\in A_{0}$, we shall see that $x^{\prime}\in A_{1}$. Indeed, $\mathbbm{1}_{A_{0}}(x^{\prime})=1$ implies $(x^{\prime},0)\in A$ i.e. $\mathbbm{1}_{A}(x^{\prime},0)=1$. As $(x^{\prime},0)\leq(x^{\prime},1)$ and $A$ is increasing, $1=\mathbbm{1}_{A}(x^{\prime},0)\leq\mathbbm{1}_{A}(x^{\prime},1)=1$. ∎
By applying conditional expectation expansion to $n$th variable of Dirichlet forms, we can readily reduce the Dirichlet form of the original function to that of the restricted functions.
Lemma 3.2.
For any $g:\{0,1\}^{n}\to\mathbb{R}$, we have
$\mathcal{E}_{p}^{n}(g,g)=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}.$ (28)
Proof.
Note that
$\begin{split}\mathcal{E}_{p}^{n}(g,g)&=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\underset{x_{n}\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},x_{n})}(g(x^{\prime},x_{n})-g(y))^{2}\right]\\ &=p\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}\right]\\ &+(1-p)\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}\right].\end{split}$
Then we decompose each sum over $V_{1}$ and $V_{0}$ to obtain
$\begin{split}\mathcal{E}_{p}^{n}(g,g)&=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+p\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\\ &+(1-p)\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}.\end{split}$
We also need to use some simple properties of the function $f(t):=\frac{\log_{p}t}{t}$.
Lemma 3.3.
The function $f(t)$ is decreasing and convex for $t\in[0,1],p\in(0,1)$. Moreover, it satisfies the identity
$f(\beta\cdot t)=\frac{\log_{p}(\beta\cdot t)}{\beta\cdot t}=\frac{1}{\beta}\cdot\frac{\log_{p}t}{t}+\frac{\log_{p}\beta}{\beta}\cdot\frac{1}{t}.$
Proof.
Note that $f^{\prime}(t)=\frac{1-\log t}{\log p\cdot t^{2}}\leq 0$ and $f^{\prime\prime}(t)=\frac{-3+2\log t}{\log p\cdot t^{3}}\geq 0$ when $t\in[0,1],p\in(0,1)$. ∎
3.2 Induction Part
We shall make several simple assumptions:
  • We may assume without loss of generality that $g\geq 0$, since replacing $g$ by its absolute value preserves the right-hand side of Eq. 8, while it only decreases the left-hand side:
    $\begin{split}\mathcal{E}_{p}(g,g)&=\underset{x\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}|g(x)-g(y)|^{2}\right]\\ &\geq\underset{x\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}(|g(x)|-|g(y)|)^{2}\right]=\mathcal{E}_{p}(|g|,|g|).\end{split}$
  • We may assume the support of $g$ is the whole set $A$. Indeed, if we denote $B:={\rm supp}(g)=\{x\in\{0,1\}^{n}:g(x)>0\}$ with $B\subset A$. Then $\mathbb{E}_{p}[g]^{2}=\left(\sum_{x\in B}g(x)\mu_{p}(x)\right)^{2}$. Hence, replacing the support of $g$ with the whole set $A$ only increases the right-hand side of Eq. 8.
It is easy to check that the theorem holds for $n=1$. Let $n\geq 2$, and suppose the statement of the theorem holds when $n$ is replaced by n1n-1. Using Lemma 3.2 to split the Dirichlet form of the original function to two restricted functions:
$\mathcal{E}_{p}^{n}(g,g)=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}.$ (29)
Here, we use the superscript $k$ with $k=n$ or n1n-1 to indicate that the Dirichlet form is the one defined on the $k$-dimensional hypercube. In the following, we adopt the same convention for other quantities, e.g., the measure $\mu_{p}$. By inductive assumption, we have
$p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})\geq\frac{\mathbb{E}_{p}[g_{1}]^{2}}{a_{1}}\log_{p}a_{1},\quad p\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\geq\frac{\mathbb{E}_{p}[g_{0}]^{2}}{a_{0}}\log_{p}a_{0}.$ (30)
Thus our goal is to show that
$\begin{split}p\mathcal{E}_{p}^{n}(g,g)&=p^{2}\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+p(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+p\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}\\ &\geq p\cdot\frac{\mathbb{E}_{p}[g_{1}]^{2}}{a_{1}}\log_{p}a_{1}+(1-p)\cdot\frac{\mathbb{E}_{p}[g_{0}]^{2}}{a_{0}}\log_{p}a_{0}+p\cdot\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}\\ &\underset{?}{\geq}\frac{\log_{p}(pa_{1}+(1-p)a_{0})}{pa_{1}+(1-p)a_{0}}\cdot\mathbb{E}_{p}[g]^{2}.\end{split}$ (31)
  • Finally, we may assume both $g_{0},g_{1}$ to be constant on $A_{0}$ and on $A_{1}\setminus A_{0}$ (and of course $g_{i}$ vanish on $A_{i}^{c}$, in particular $g_{0}$ is 0 on $A_{1}\setminus A_{0}$). Indeed, if we replace $g_{j}$ with $\mathbb{E}_{p}[g_{j}]$ for $j=0,1$, then
    $\begin{split}&\left\langle\mathbbm{1}_{A_{j}},\mathbb{E}_{x\sim\mu_{p}}[g_{j}]\right\rangle_{p}=\underset{x^{\prime}\sim\mu_{p}}{\mathbb{E}}[g_{j}(x^{\prime})\cdot\mathbbm{1}_{A_{j}}]\cdot a_{j}\leq\underset{x^{\prime}\sim\mu_{p}}{\mathbb{E}}[g_{j}(x^{\prime})\cdot\mathbbm{1}_{A_{j}}],\\ &\|\mathbb{E}_{x\sim\mu_{p}}[g_{1}]-\mathbb{E}_{x\sim\mu_{p}}[g_{0}]\|_{2,\mu_{p}}^{2}\leq\underset{x\sim\mu_{p}}{\mathbb{E}}[\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}]=\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2},\end{split}$
    where the first inequality is due to $a_{j}=\mu_{p}^{n-1}(A_{j})\leq 1$ and the second inequality is due to Jensen’s inequality. Hence replacing $g_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of Eq. 31.
Now let $\alpha$ be the value of $g_{1}$ on $A_{0}$ and $\gamma$ be the value of $g_{1}$ on $A_{1}\setminus A_{0}$. Let $\beta$ be the value of $g_{0}$ on $A_{0}$. Note that
$\begin{split}&s_{1}:=\mathbb{E}_{p}[g_{1}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\alpha a_{0}+\gamma(a_{1}-a_{0}),\\ &s_{0}:=\mathbb{E}_{p}[g_{0}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)=\beta a_{0},\\ &\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &\quad\ =\sum\limits_{x^{\prime}\in A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})+\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})\\ &\quad\ =(\alpha-\beta)^{2}\cdot a_{0}+(a_{1}-a_{0})\cdot\gamma^{2}.\end{split}$
Then
[…]
With the new notation, the inequality to be verified for the induction step is:
$\begin{split}p\cdot\frac{\log_{p}a_{1}}{a_{1}}\cdot s_{1}^{2}&+(1-p)\cdot\frac{\log_{p}a_{0}}{a_{0}}\cdot s_{0}^{2}+p\cdot(\alpha-\beta)^{2}\cdot a_{0}+p\cdot(a_{1}-a_{0})\cdot\gamma^{2}\\ &\geq\frac{\log_{p}(pa_{1}+(1-p)a_{0})}{pa_{1}+(1-p)a_{0}}\cdot(ps_{1}+(1-p)s_{0})^{2}.\end{split}$ (32)
Let $f(t):=\frac{\log_{p}t}{t}$ with $t\in[0,1]$. Then our goal is to prove that
$\begin{split}ps_{1}^{2}f(a_{1})+(1-p)s_{0}^{2}f(a_{0})&+pa_{0}(\alpha-\beta)^{2}+p(a_{1}-a_{0})\gamma^{2}\\ &\geq(ps_{1}+(1-p)s_{0})^{2}f(pa_{1}+(1-p)a_{0})).\end{split}$ (33)
Expressing $\beta$ and $\gamma$ as functions of $s_{i},a_{i}$ and of $\alpha$, LHS of the above is a quadratic in $\alpha$ with coefficients depending on $s_{i}$ and $a_{i}$: LHS $A\alpha^{2}+B\alpha+C$, where
$\begin{split}&A:=p\left(a_{0}+\frac{a_{0}^{2}}{a_{1}-a_{0}}\right)\\ &B:=-2p\left(s_{0}+\frac{a_{0}s_{1}}{a_{1}-a_{0}}\right)\\ &C:=ps_{1}^{2}f(a_{1})+(1-p)s_{0}^{2}f(a_{0})+p\frac{s_{0}^{2}}{a_{0}}+p\frac{s_{1}^{2}}{a_{1}-a_{0}}.\end{split}$
Minimizing LHS in $\alpha$ (when $\alpha=-\frac{B}{2A}=\frac{s_{0}a_{1}-s_{0}a_{0}+a_{0}s_{1}}{a_{1}a_{0}}$), we arrive, after some simple calculations, to the following inequality we need to verify:
$ps_{1}^{2}f(a_{1})+(1-p)s_{0}^{2}f(a_{0})+p\frac{(s_{0}-s_{1})^{2}}{a_{1}}\geq(ps_{1}+(1-p)s_{0})^{2}f(pa_{1}+(1-p)a_{0})).$
Next, let $T=\frac{s_{0}}{s_{1}}$. The above inequality transforms to a quadratic inequality in $T$:
$pf(a_{1})+(1-p)f(a_{0})T^{2}+p\frac{(1-T)^{2}}{a_{1}}\geq(p+(1-p)T)^{2}f(pa_{1}+(1-p)a_{0})).$ (34)
We need to check $h(T):=A^{\prime}T^{2}+B^{\prime}T+C^{\prime}\geq 0$ with the coefficients
$\begin{split}&A^{\prime}:=(1-p)f(a_{0})+\frac{p}{a_{1}}-(1-p)^{2}f(pa_{1}+(1-p)a_{0}))\\ &B^{\prime}:=-\frac{2p}{a_{1}}-2p(1-p)f(pa_{1}+(1-p)a_{0}))\\ &C^{\prime}:=pf(a_{1})+\frac{p}{a_{1}}-p^{2}f(pa_{1}+(1-p)a_{0})).\end{split}$
We shall verify $A^{\prime}\geq 0$ and $D=B^{\prime 2}-4A^{\prime}C^{\prime}\leq 0$, which will conclude the proof.
Lemma 3.4.
$A^{\prime}:=(1-p)f(a_{0})+\frac{p}{a_{1}}-(1-p)^{2}f(pa_{1}+(1-p)a_{0}))\geq 0$.
Proof.
Note that
$\begin{split}A^{\prime}&=(1-p)f(a_{0})+\frac{p}{a_{1}}-(1-p)^{2}f(pa_{1}+(1-p)a_{0}))\\ &=\frac{(1-p)a_{1}f(a_{0})+p-(1-p)^{2}a_{1}f(pa_{1}+(1-p)a_{0}))}{a_{1}}.\end{split}$
By Lemma 3.3, $f(pa_{1}+(1-p)a_{0}))\leq f(a_{0})$ and hence,
$(1-p)^{2}a_{1}f(pa_{1}+(1-p)a_{0}))\geq-(1-p)^{2}a_{1}f(a_{0})\geq-(1-p)a_{1}f(a_{0}).$
Then
$A^{\prime}\geq\frac{(1-p)a_{1}f(a_{0})+p-(1-p)a_{1}f(a_{0})}{a_{1}}=\frac{p}{a_{1}}\geq 0.$
It remains to verify the inequality $4A^{\prime}C^{\prime}\geq B^{\prime 2}$, which reduces to a two-point inequality:
$\begin{split}&pf(a_{1})+(1-p)f(a_{0})+(1-p)a_{1}f(a_{0})f(a_{1})\\ &\geq(p(1-p)a_{1}f(a_{0})+(1-p)^{2}a_{1}f(a_{1})+1)f(pa_{1}+(1-p)a_{0}).\end{split}$ (35)
3.3 Two-point Inequality Part
Renaming the variables $x=a_{0}$ and $y=a_{1}$, and recalling the constraints on $a_{0}$ and $a_{1}$, we need to prove Eq. 35 for $0<x<y\leq 1$.
$\begin{split}&p\cdot f(y)+(1-p)\cdot f(x)+(1-p)\cdot yf(x)f(y)\\ &\geq(p(1-p)yf(x)+(1-p)^{2}yf(y)+1)\cdot f(py+(1-p)x).\end{split}$ (36)
Rearranging, this is easily seen to be equivalent to
$\Delta_{p}(x,y)\geq\frac{p(1-p)^{2}y(f(x)-f(y))^{2}}{p(1-p)yf(x)+(1-p)^{2}yf(y)+1},$ (37)
where
$\Delta_{p}(x,y):=pf(y)+(1-p)f(x)-f(py+(1-p)x).$
Note that $\Delta_{p}(x,y)\geq 0$ since $f(t):=\frac{\log_{p}t}{t}$ is convex.
We now substitute $x=ty$ in $\Delta_{p}(x,y)$, with $0<t<1$, and expand it using Lemma 3.3. We have
$\begin{split}&\Delta_{p}(x,y)=\Delta_{p}(ty,y)=pf(y)+(1-p)f(t\cdot y)-f((p+(1-p)t)\cdot y)\\ &=pf(y)+\frac{1}{t}\cdot(1-p)f(y)+\frac{1}{y}\cdot(1-p)f(t)-\frac{f(y)}{p+(1-p)t}-\frac{f(p+(1-p)t)}{y}\\ &=\frac{p(1-p)(t-1)^{2}}{t(p+(1-p)t)}\cdot f(y)+\frac{1}{y}\cdot\left(\frac{(1-p)\log_{p}t}{t}-\frac{\log_{p}(p+(1-p)t)}{p+(1-p)t}\right).\end{split}$
As to the right-hand side of Eq. 37, we have
$\begin{split}{\rm RHS}(x,y)&={\rm RHS}(ty,y)=\frac{p(1-p)^{2}y[(\frac{1}{t}-1)f(y)+\frac{1}{y}\cdot\frac{\log_{p}t}{t}]^{2}}{[\frac{p(1-p)}{t}+(1-p)^{2}]yf(y)+p(1-p)\cdot\frac{\log_{p}t}{t}+1}\\ &=\frac{y[(\frac{1}{t}-1)f(y)+\frac{1}{y}\cdot\frac{\log_{p}t}{t}]^{2}}{[\frac{1}{t(1-p)}+\frac{1}{p}]yf(y)+\frac{\log_{p}t}{(1-p)t}+\frac{1}{p(1-p)^{2}}}.\end{split}$
Taking $z:=yf(y)$,
$\Delta_{p}(x,y)\geq{\rm RHS}\Leftrightarrow y\cdot\Delta_{p}(x,y)\geq y\cdot{\rm RHS}\Leftrightarrow Az+B\geq\frac{(Cz+D)^{2}}{Ez+F},$
where
$z=yf(y)=\log_{p}(y)>0\ \text{as $0<y<1$},$
and $A,B,\dots,F$ depend only on $t$ and $p$. Specifically,
$\begin{cases}A=\frac{p(1-p)(t-1)^{2}}{t(p+(1-p)t)},\\ B=\frac{(1-p)\log_{p}t}{t}-\frac{\log_{p}(p+(1-p)t)}{p+(1-p)t},\\ C=\frac{1-t}{t},\\ D=\frac{\log_{p}t}{t},\\ E=\frac{1}{t(1-p)}+\frac{1}{p}\geq 0,\\ F=\frac{\log_{p}t}{(1-p)t}+\frac{1}{p(1-p)^{2}}\geq 0.\end{cases}$
Thus we need to show that
$(Az+B)(Ez+F)\geq(Cz+D)^{2}.$
Observe that
$AE=\frac{(t-1)^{2}}{t^{2}}=C^{2}.$
Therefore, this reduces to a linear inequality in $z$:
$(AF+BE-2CD)\cdot z\geq D^{2}-BF.$
This holds for all non-negative $z$ if and only if $AF+BE\geq 2CD$ and $BF\geq D^{2}$.
Lemma 3.5.
For any $t\in(0,1)$ and $p\in(0,1)$ we have $AF+BE\geq 2CD$.
Proof.
Simplifying and rearranging terms, the inequality reduces to verifying that
$F(t,p):=(1-p)t\log_{p}t+p(t-1)^{2}-(p+(1-p)t)\log_{p}(p+(1-p)t)\geq 0.$
We compute the derivatives of $F(t,p)$ with respect to $t$:
$\begin{split}&\frac{\partial}{\partial t}F(t,p)=\frac{1-p}{\ln(p)}\cdot\ln\left(\frac{t}{p+(1-p)t}\right)+2p\cdot(t-1),\\ &\frac{\partial^{2}}{\partial t^{2}}F(t,p)=\frac{p(1-p)}{\ln(p)}\cdot\frac{1}{t(p+(1-p)t)}+2p,\\ &\frac{\partial^{3}}{\partial t^{3}}F(t,p)=\frac{p(1-p)}{\ln(1/p)}\cdot\frac{p+2(1-p)t}{t^{2}(p+(1-p)t)^{2}}\geq 0.\end{split}$
Hence $\frac{\partial}{\partial t}F(t,p)$ is a convex function of $t$, implying that it can vanish at most twice on the interval $t\in(0,1)$. Note that $\frac{\partial}{\partial t}F(t,p)|_{t\to 0}>0$, $\frac{\partial}{\partial t}F(t,p)|_{t=1}=0$. Therefore, there exists $t_{0}:=t_{0}(p)\in(0,1)$ such that $\frac{\partial}{\partial t}F(t,p)|_{t=t_{0}}=0$ and $\frac{\partial}{\partial t}F(t,p)\geq 0$ for $t\in(0,t_{0}]$ and $F(t,p)\leq 0$ for $t\in(t_{0},1)$. Since $F(0,p)=F(1,p)=0$, the function $F(t,p)$ first increases from 0 at $t=0$ and then decreases back to 0 at $t=1$. Consequently, $F(t,p)\geq 0$ for all $t\in(0,1)$, which establishes the desired inequality. ∎
Lemma 3.6.
For any $t\in(0,1)$ and $p\in(0,1)$ we have $BF\geq D^{2}$.
Proof.
Simplifying and rearranging terms, this inequality reduces to show that
$\begin{split}&G(t,p):=(1-p)(p+(1-p)t)\log_{p}t\\ &-p(1-p)\log_{p}(t)\cdot\log_{p}(p+(1-p)t)-t\cdot\log_{p}(p+(1-p)t)\geq 0.\end{split}$
Clearly, $G(0,p)=G(1,p)=0$. We aim to show that $G(t,p)$ increases from 0 at $t=0$, attains its maximum at some interior point, and then decreases back to 0 at $t=1$; hence $G(t,p)\geq 0$ for all $t\in(0,1)$.
We begin by computing the derivative:
$\begin{split}\frac{\partial}{\partial t}G(t,p)&=\frac{1}{\ln(p)}\left(\frac{p(1-p)}{t}+(1-p)^{2}\ln(t)+(1-p)^{2}\right)\\ &-\frac{1}{\ln(p)}\left(\frac{p(1-p)}{\ln(p)}\cdot\frac{\ln(p+(1-p)t)}{t}\right)\\ &-\frac{1}{\ln(p)}\left(\frac{p(1-p)^{2}}{\ln(p)}\cdot\frac{\ln(t)}{p+(1-p)t}+\ln(p+(1-p)t)+\frac{(1-p)t}{p+(1-p)t}\right).\end{split}$
Define $H(t,p):=t(p+(1-p)t)\cdot\frac{\partial}{\partial t}G(t,p)$. To analyze its sign, we consider successive derivatives of $H$.
$\begin{split}&\frac{\partial}{\partial t}H(t,p)=\frac{1}{\ln(p)^{2}}\left[-(p-1)^{2}p\left(2+\ln(t)+\ln(p+t-pt)\right)\right]\\ &+\frac{1}{\ln(p)}\left[(-2t+p(2t-1)\ln(p+t-pt)\right]\\ &-\frac{1}{\ln(p)}\left[3(p-1)p(1+p(t-1)-2t)+(p-1)^{2}(-2t+p(2t-1))\ln(t)\right];\\ &\frac{\partial^{2}}{\partial t^{2}}H(t,p)=\frac{1}{\ln(p)^{2}}\left(-(p-1)^{2}p\left(\frac{1}{t}+\frac{1-p}{p+t-pt}\right)\right)\\ &+\frac{1}{\ln(p)}\left(-3(-2+p)(p-1)p-\frac{(p-1)^{2}(-2t+p(2t-1))}{t}\right)\\ &+\frac{1}{\ln(p)}\left(\frac{(1-p)(-2t+p(2t-1))}{p+t-pt}-2(p-1)^{3}\ln(t)+2(p-1)\ln(p+t-pt)\right);\\ &\frac{\partial^{3}}{\partial t^{3}}H(t,p)=-\frac{(p-1)^{2}p}{t^{2}(p+t-pt)^{2}\ln(p)^{2}}\cdot h(t,p),\end{split}$
where $h(t,p):=-2t^{2}+2pt(-1+2t)+p^{2}(-1+2t-2t^{2})+(-6p(-1+t)t^{2}+4t^{3}+p^{2}(t-1)^{2}(1+2t))\ln(p)$. Thus it suffices to consider the sign of function $h(t,p)$ for $t\in(0,1)$ and $p\in(0,1)$. Note that
$\begin{split}&\frac{\partial}{\partial t}h(t,p)=-2(p-1)(-2t+p(-1+2t))+6(p-2)(p(t-1)-t)t\ln(p),\\ &\frac{\partial^{2}}{\partial t^{2}}h(t,p)=-4(p-1)^{2}+6(p-2)(-2t+p(-1+2t))\ln(p),\\ &\frac{\partial^{3}}{\partial t^{3}}h(t,p)=12(p-2)(p-1)\ln(p)<0.\end{split}$
Hence $\frac{\partial^{2}}{\partial t^{2}}h(t,p)$ is strictly decreasing on $(0,1)$. Evaluating at $t\to 0$, we obtain
$\lim_{t\to 0}\frac{\partial^{2}}{\partial t^{2}}h(t,p)=-4(p-1)^{2}-6(p-2)p\ln(p).$
Write $f(p):=-4(p-1)^{2}-6(p-2)p\ln(p)$. Note that f(0)=4,f(1)=0f(0)=-4,f(1)=0, and
$\begin{split}&f^{\prime}(p)=-2(-10+7p+6(p-1)\ln(p)),\\ &f^{\prime\prime}(p)=-26+\frac{12}{p}-12\ln(p),\\ &f^{\prime\prime\prime}(p)=-\frac{12(1+p)}{p^{2}}<0.\end{split}$
Therefore $f^{\prime\prime}(p)$ has a unique zero $p_{0}\in(0,1)$ since $f^{\prime\prime}(0)\to\infty$ and $f^{\prime\prime}(1)=-14$. It follows that $f^{\prime}(p)$ increases on $(0,p_{0})$ and decreases on $(p_{0},1)$. As $f^{\prime}(0)\to$-$\infty$ and $f^{\prime}(1)=6$, we deduce that $f(p)$ first decreases on $(0,p_{0})$ and then increases on $(p_{0},1)$, so $f(p)=\lim_{t\to 0}\frac{\partial^{2}}{\partial t^{2}}h(t,p)\leq 0$ for all $p\in(0,1)$. Consequently, $\frac{\partial^{2}}{\partial t^{2}}h(t,p)$ is negative, and since $\frac{\partial}{\partial t}h(t,p)\leq\frac{\partial}{\partial t}h(t,p)|_{t=0}=2p(p-1)<0$, we have $h(t,p)<0$ throughout $(0,1)$. Thus, $\frac{\partial^{3}}{\partial t^{3}}H(t,p)>0$.
Next, note that $\lim_{t\to 0}\frac{\partial^{2}}{\partial t^{2}}H(t,p)\to$-$\infty$ and $\frac{\partial^{2}}{\partial t^{2}}H(t,p)|_{t=1}>0$ with $\frac{\partial^{2}}{\partial t^{2}}H(t,p)$ increasing on $t\in(0,1)$. Hence it has a unique zero $t_{0}\in(0,1)$, and $\frac{\partial}{\partial t}H(t,p)$ decreases on $(0,t_{0})$ then increases on $(t_{0},1)$. Since $\frac{\partial}{\partial t}H(t,p)|_{t=0},\frac{\partial}{\partial t}H(t,p)|_{t=1}>0$ and $\frac{\partial}{\partial t}H(t,p)|_{t=t_{0}}<0$, there exists $0<t_{1}<t_{2}<1$ such that $H(t,p)$ first increases on $(0,t_{1})$, then decreases on $t_{1},t_{2}]$ and finally increases again on $(t_{2},1)$. Because $H(0,p)=H(1,p)=0$, there exists $t^{\prime}\in[0,1]$ such that $H(t,p)>0$ for $t\in(0,t^{\prime})$ and $H(t,p)\leq 0$ for $t\in[t^{\prime},1)$. It implies that $G(t,p)$ increases on $t\in(0,t^{\prime})$ and then deceases in $t\in[t^{\prime},1)$. Since $G(0,p)=G(1,p)=0,$ we conclude that $G(t,p)\geq 0$ for all $t\in[0,1]$.
4 Proof of Theorem 1.7
We now show how to derive Theorem 1.7 from the $p$-biased discrete functional inequality Eq. 8 established in Theorem 1.6.
For $g:\{0,1\}^{n}\to\mathbb{R}$, recall the $p$-biased Dirichlet form
$\mathcal{E}_{p}(g,g):=\sum_{x}\mu_{p}(x)\sum_{y\sim x}\left(g(x)-g(y)\right)^{2}.$
Let $L_{p}:=D-W$ be the (external) weighted Laplacian on $A$. Then the standard edge-vertex double counting gives
$\mathcal{E}_{p}(g,g)=\sum_{\{x,y\}\subseteq A}\omega_{p}(x,y)\bigl(g(x)-g(y)\bigr)^{2}=\vec{g}^{\top}L_{p}\vec{g}.$ (38)
Indeed, grouping the two directed contributions $(x,y)$ and $(y,x)$ across each undirected edge $\{x,y\}$ yields
$\begin{split}\mu_{p}(x)(g(x)-g(y))^{2}+\mu_{p}(y)(g(x)-g(y))^{2}&=(\mu_{p}(x)+\mu_{p}(y))(g(x)-g(y))^{2}\\ &=\omega_{p}(x,y)(g(x)-g(y))^{2};\end{split}$
expanding the rightmost equality is the usual identity $\sum_{\{x,y\}}\omega_{p}(x,y)(g(x)-g(y))^{2}=\vec{g}^{\top}(D-W)\vec{g}$.
We may assume without loss of generality that $g\geq 0$. Indeed, since $||g(x)|-|g(y)||\leq|g(x)-g(y)|$, replacing $g$ by $|g|$ can only decrease the Dirichlet energy, while $\mathbb{E}_{\mu_{p}}[|g|]\geq|\mathbb{E}_{\mu_{p}}[g]|$ ensures that the right-hand side of Eq. 8 does not decrease. Therefore, Theorem 1.6 gives
$p\cdot\mathcal{E}_{p}(g,g)\geq\frac{(\mathbb{E}_{\mu_{p}}g)^{2}}{\mu_{p}(A)}\cdot\log_{p}(\mu_{p}(A)).$ (39)
Let $\vec{\mu}=(\mu_{p}(x))_{x\in A}$, so that $\mathbb{E}_{\mu_{p}}[g]=\left\langle\vec{\mu},\vec{g}\right\rangle$. Inserting (38) into (39) gives
$\vec{g}^{\top}L_{p}\vec{g}\ \geq\rho\cdot(\vec{\mu}^{\top}\vec{g})^{2},\quad\rho:=\frac{\log_{p}\left(\mu_{p}(A)\right)}{p\mu_{p}(A)}.$
Equivalently,
$L_{p}\succeq\rho\cdot\vec{\mu}\vec{\mu}^{\top}\Rightarrow\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\leq\frac{1}{\rho}=\frac{p\mu_{p}(A)}{\log_{p}\big(\mu_{p}(A)\big)}.$ (40)
Since $L_{p}=D(I-P_{A})$, we have
$L_{p}^{-1}=(I-P_{A})^{-1}D^{-1}=\sum_{k\geq 0}P_{A}^{k}D^{-1}.$ (41)
Let $\vec{\nu}$ be the initial distribution $\nu(x)=\mu_{p}(x)/\mu_{p}(A)$ on $A$ and set $\vec{\phi}:=D^{-1}\vec{\mu}$, i.e. $\phi(x)=\mu_{p}(x)/d_{p}(x)$. Let $(X_{t})_{t\geq 0}$ be the Markov chain with transition $P_{A}$, killed upon exiting $A$, and let $\tau:=\min\{t\geq 1:X_{t}\notin A\}$. By the tower property and the Markov property, for every $k\geq 0$ and every $f:\{0,1\}^{n}\to\mathbb{R}$ supported on $A$,
$(P_{A}^{k}f)(x)\ =\ \mathbb{E}_{x}\left[f(X_{k})\mathbbm{1}_{\{\tau>k\}}\right],$ (42)
where $\mathbb{E}_{x}$ is the expectation taken with respect to a $p$-biased random walk starting at a given point $x$. Applying (42) with $f=\phi$ and averaging over $\vec{\nu}$,
$\vec{\nu}^{\top}P_{A}^{k}\vec{\phi}=\mathbb{E}_{\nu}\left[\phi(X_{k})\mathbbm{1}_{\{\tau>k\}}\right].$
Using (41) we obtain
$\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}=\sum_{k\geq 0}\vec{\mu}^{\top}P_{A}^{k}D^{-1}\vec{\mu}=\mu_{p}(A)\sum_{k\geq 0}\vec{\nu}^{\top}P_{A}^{k}\vec{\phi}=\mu_{p}(A)\sum_{k\geq 0}\mathbb{E}_{\nu}\left[\phi(X_{k})\mathbbm{1}_{\{\tau>k\}}\right].$ (43)
Combining with Eq. 40 proves
$\tilde{T}(A,p)=\sum_{k\geq 0}\mathbb{E}_{\nu}\left[\phi(X_{k})\mathbbm{1}_{\{\tau>k\}}\right]=\frac{1}{\mu_{p}(A)}\cdot\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\leq\frac{p}{\log_{p}\big(\mu_{p}(A)\big)},$
where the equality case is exactly the extremal case of Theorem 1.6, namely that $A$ is an increasing subcube.
For $x$ with Hamming weight $t=|x|$, a direct calculation gives
$\phi(x)=\frac{\mu_{p}(x)}{d_{p}(x)}\ =\ \frac{1}{\,n+(n-t)\frac{p}{1-p}+t\frac{1-p}{p}\,}.$ (44)
Hence for all $x\in A$,
$\frac{\min\{p,1-p\}}{n}\ \leq\ \phi(x)\ \leq\ \frac{\max\{p,1-p\}}{n}.$ (45)
From (43) and (45) we get
$\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\geq\frac{\min\{p,1-p\}}{n}\mu_{p}(A)\sum_{k\geq 0}\mathbb{P}_{\nu}(\tau>k)=\frac{\min\{p,1-p\}}{n}\mu_{p}(A)\mathbb{E}_{\nu}[Y].$ (46)
Combining (40) and (46) yields
$\frac{\min\{p,1-p\}}{n}\mu_{p}(A)\mathbb{E}_{\nu}[Y]\leq\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\leq\frac{p\mu_{p}(A)}{\log_{p}\big(\mu_{p}(A)\big)},$
hence
$T(A,p)=\mathbb{E}_{\nu}[Y]\leq\frac{p}{\min\{p,1-p\}}\cdot\frac{n}{\log_{p}\big(\mu_{p}(A)\big)}.$
Equality case. Equality in the final bound requires equality in each step where an inequality was used:
(a) Equality in (40) is equivalent to equality in the functional inequality (39). By the equality characterization of Theorem 1.6, this holds if and only if $A$ is an increasing subcube.
(b) Equality in (46) forces $\phi(X_{k})$ to be almost surely constant along all surviving paths. From (44) one sees that $\phi$ depends linearly on $|x|$ and is constant on AA only when $p=\tfrac{1}{2}$ (in which case $\phi\equiv 1/(2n)$). Thus the equality case occurs exactly when $p=\tfrac{1}{2}$ and $A$ is an increasing subcube.
5 Proof of mean first exit time for increasing subcubes
In this section, we verify the mean survival time of random walks from increasing subcubes (1). For $x\in A$, write $t(x):=\sum_{i=k+1}^{n}x_{i}$ for the number of free $1$’s. By permutation symmetry among the free coordinates, the conditional expectation of the exit time only depends on $t(x)$. Set
$m(t):=\mathbb{E}[Y\mid t(X_{0})=t],\ t=0,1,\dots,n-k.$
From a state with value $t$, exactly three mutually exclusive one-step outcomes can occur:
  • (exit) flip one of the first $k$ fixed $1$’s to 0 — this exits $A$;
  • (down) flip a free $1$ to 0, so $t\mapsto t-1$;
  • (up) flip a free 0 to $1$, so $t\mapsto t+1$.
Counting the three kinds of neighbors (there are $k$ exit-edges, $t$ down-edges, and nktn-k-t up-edges), we obtain
$\frac{d_{p}(x)}{\mu_{p}(x)}=\frac{k+t}{p}+\frac{n-k-t}{1-p}.$
Define
$D(t):=p(1-p)\frac{d_{p}(x)}{\mu_{p}(x)}=(1-p)(k+t)+p\bigl((n-k)-t\bigr),$
and hence the three one-step probabilities from level $t$ are
$e(t)=\frac{k(1-p)}{D(t)},\ d(t)=\frac{t(1-p)}{D(t)},\ u(t)=\frac{(n-k-t)p}{D(t)},\ e(t)+d(t)+u(t)=1.$
By the Markov property (first-step analysis), conditioning on the three one-step outcomes gives, for every $t\in\{0,\dots,n-k\}$,
$m(t)=1+d(t)\cdot m(t-1)+u(t)\cdot m(t+1),$ (47)
with d(0)=u(nk)=0d(0)=u(n-k)=0. Now let us compute the equation
$D(t)m(t)=D(t)+t(1-p)\cdot m(t-1)+(n-k-t)p\cdot m(t+1).$
Substituting $m(t)=A+Bt$, expanding both sides and matching coefficients of $t^{2},t^{1}$ and the constant term yields
$B=\frac{1-2p}{1+(1-p)k},\quad A=1+\frac{p(n-k)}{(1-p)k}(1+B).$
Under the starting law $\nu=\mu_{p}(\cdot)/\mu_{p}(A)$, the free coordinates are i.i.d. ${\rm Ber}(p)$, so $T_{0}:=t(X_{0})\sim{\rm Bin}(n-k,p)$ and $\mathbb{E}[T_{0}]=p(n-k)$. Hence
$T(C,p)=\mathbb{E}_{\nu}[Y]=\mathbb{E}[m(T_{0})]=A+B\cdot\mathbb{E}[T_{0}]=1+2p\left(\frac{n}{k}-1\right).$
This completes the proof.
6 Conclusion and open problems
A natural next target beyond our $O(n^{2})$ mixing bound for constant-density increasing sets is the full Ding–Mossel conjecture asking for an $O(n\log n)$ mixing time for the censored walk when $\mu(A)\geq\varepsilon>0$ ([DM2014, Question 1.1]). As observed in [fei2025spectral], a route to such an $O(n\log n)$ bound is to establish an appropriate (standard or modified) log-Sobolev inequality for the censored chain on $A$. Following the general relations between log-Sobolev and mixing time bound (e.g.,[BT2006logsobolev, DS1996logsobolev]), it is natural to ask what the correct dependence on the density $\mu(A)$ should be in any valid log–Sobolev (or modified log-Sobolev) inequality for increasing sets.
Our inductive two-point machinery is well suited for quadratic ($L^{2}$) control: it converts high-dimensional Dirichlet forms and variances into lower-dimensional analogues plus a finite list of analytics inequalities (e.g., two-point inequalities) that one can verify exactly. However, in log-Sobolev inequalities, the relative entropy is a nonlinear, non-quadratic functional—this introduces challenges when applying the induction method to control it. Hence, to establish log-Sobolev inequalities in non-product spaces, further developing the induction method or looking for new methods is needed.